OpenStudy (anonymous):

evaluate the expression for a=-1 and b=5 (ab)2

OpenStudy (anonymous):

replace a in (ab)2 with -1 and b with 5

OpenStudy (anonymous):

-10 is what I got

OpenStudy (anonymous):

do I add or multiply

OpenStudy (anonymous):

is that (a*b)^2 or a*b*2

OpenStudy (anonymous):

if I multiply I got 10 if I add I got 8

OpenStudy (anonymous):

what is ^

OpenStudy (anonymous):

what does it mean

OpenStudy (anonymous):

^ means squared.

OpenStudy (anonymous):

(a*b*2) (-1*5*2) = -5*2 = -10

OpenStudy (anonymous):

(a*b)^2 (-1*5)^2 (-5)^2 = 25

OpenStudy (anonymous):

ok that other person got me confused I knew it was -10 or 8

OpenStudy (anonymous):

so is it 10 or 25

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

it's -10

OpenStudy (anonymous):

they just got me confused I knew I was on the right track